Four 20 Ω resistors are connected in series and the combination is connected to a 20 V emf device. The current in any one of the resistors is:
Answer #1

Equivalent resistance for series combination of \(n\) resistances is given by

\(R_{e q}=R_{1}+R_{2}+R_{3} \ldots \ldots \ldots \ldots \ldots . .+R_{n}\)

In our case we have \(\mathrm{n}=4\) and each resistance is of \(20 \mathrm{ohm} \mathrm{so}\)

\(R_{e q}=20+20+20+20=80 o h m\)

Given Emf of cell is, \(E=20 \mathrm{~V}\)

Using ohm's law current in circuit is

\(I=\frac{E}{R_{e q}}\)

\(I=\frac{20}{80}\)

\(I=0.25 A\)